54.000/12=4500
8%=0.08
0.08*4500=360
360-100=260
How do I write this as an inequality? Car payments. As a rule of thumb, debt payments (other than mortgages) should be less than 8% of a consumer’s monthly gross income. Oliver makes $54,000 a year and has a $100 student-loan payment every month. What size car payment can he afford?
Divide 54,000 by 12 which will =4,500 Take 8%turn it into a decimal 0.08and multiply 4500=$360 then subtract the student loan $360-$100=$260
Oliver can afford a $260 car payment.

Answer by Alan3354(69443)

54.000/12=4500
8%=0.08
0.08*4500=360
360-100=260
How do I write this as an inequality? Car payments. As a rule of thumb, debt payments (other than mortgages) should be less than 8% of a consumer’s monthly gross income. Oliver makes $54,000 a year and has a $100 student-loan payment every month. What size car payment can he afford?
Divide 54,000 by 12 which will =4,500 Take 8%turn it into a decimal 0.08and multiply 4500=$360 then subtract the student loan $360-$100=$260
Oliver can afford a $260 car payment.
-----------
You have it right, unless Oliver has other debts he didn't disclose.